News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 10 July 2009 9.00am EDT

On July 7 the OBC grating "move failed" flag was unexpectedly set following a nominal retraction of the HETG. Initial engineering assessment indicates that the cause was a change in state of the HETG retract limit switch that has been unused since Aug 2000 following anomalous behavior. Since then, grating operations have been conducted using a software patch to avoid using the anomalous switch. Given the new behavior a flight software K-constant patch will be developed to ignore all states of the switch.

Following the anomaly, loads were planned to remove the remaining grating observations for the week. Observations of 4U1724-307, Mkn 421 and SN1987A will be replanned in a later load. There was no impact to science due to the anomaly.

A real-time procedure was completed on Jul 7 to update the on-board ephemeris.

Of note this week is the Supernova Remnants and Pulsar Wind Nebulae in the Chandra Era science workshop being held in Boston.

The schedule of targets for the next week is shown below.

------------------------------------------
PSRJ1622-49 ACIS-I Jul 10 NGC 6342 ACIS-S Abel l2276 ACIS-I 4C73.08 ACIS-I Jul 11 Radiation Belts SGR0418+5729 HRC-I Jul 12 MKN 6 ACIS-S NGC 4051 ACIS-S Jul 13 SDSSJ1644+4157 ACIS-S HD201601 ACIS-I G1.9+0.3 ACIS-S Radiation Belts Jul 14 1RXSJ212700.3+101108 ACIS-S XTEJ1701-462 ACIS-S Terzan 5 ACIS-S Jul 15 F02021-2103 ACIS-S Jul 16 NGC 70 ACIS-I Mrk 209 ACIS-S AXJ154905-5420 ACIS-S Radiation Belts Jul 17 Vela PWN ACIS-S N49 ACIS-S Jul 18 G1.9+0.3 ACIS-S IGRJ11321-5311 ACIS-I PNDS1 ACIS-S Jul 19 NGC2305 ACIS-S

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports